A Hotel in Ville has 1,000 rooms , and the hotel charge customer $ 100 . All the hotel rooms are always rented in any given day . a . To raise revenue , the governor decides to charge hotels a tax of $ 10 per rented room . After the tax is imposed , the price for hotel rooms rises to $ 108 , and the number of rooms rented falls to 900. Calculate the amount of revenue this tax raises for Ville and the deadweight loss of the tax . b . The governor now doubles the tax to $ 20 . The price rises to $ 116 , and the number of rooms rented falls to 800. Calculate tax revenue and deadweight loss with this larger tax . Are they double , more than double , or less than double your answers in part ( a ) ? Explain .

1
Expert's answer
2022-03-13T18:53:14-0400

a.) Revenue

"Revenue=New Price\times No of rooms=108\times900=97,200"

Dead weight loss=100,000-97,200=$2800

b) New revenue

"Revenue=116\times800=92,000"

Dead weight 100000-92800=$7,200